# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

### v2

@@ 5.3 Document-oriented VQA
To validate the strong capability of our model in other document understanding scenarios, we conduct experiments on several public benchmarks and compare PDF-WuKong with other representative models. First, we evaluate the performance of PDF-WuKong on single-page document datasets [35, 37, 40]. As shown in Tab. 6, our model achieves comparable performance on single-page document understanding. This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question-answering tasks.
In addition, we assess the performance of traditional specialized models and MLLMs on two existing multi-page document QA datasets. The results shown in Tab. 7 indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document MLLM DocOwl2 [25]. Notably, on complex multi-page document datasets like DUDE [17], PDF-WuKong outperforms GPT-4V [44]. This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
Furthermore, we conduct zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18]. As shown in Tab. 8, our model uses the fewest parameters yet achieves the second-best performance. Although InternVL-V1-5-RAG surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model. Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, while that of other models significantly decreases. At a context length of 64K, PDF-WuKong even achieves the best performance, demonstrating its robustness in handling long-context multimodal inputs.
@@
Qwen-VL [57] 65.1 65.7 35.4
Monkey [58] 66.5 65.1 36.1
Text-Monkey [10] 73.0 66.9 28.6
MiniCPM-V-2.5 [60] 84.8 - -
Vary-base [12] 76.3 66.1 -
TextHawk [62] 76.4 66.6 50.6
IXC2-4KHD-16 [24] 84.9 80.1 60.8
DocOwl 2 [25] 80.7 70.0 46.4
CREAM [26] 79.4 - 53.6
@@
LayoutLMv3 [54] 55.1 20.3
Longformer [55] 55.1 27.1
BigBird [56] 58.5 26.3
Hi-VT5 [46] 61.8 35.7
DocFormerv2 [63] 76.4 48.4
GRAM [16] 83.0 53.4
GPT-4V (2024-06) [44] - 53.9
Idefics3-8B [64] 67.2 38.7
DocOwl2 [25] 69.4 46.7
CREAM [26] 65.3 52.5
@@
Emu2-Chat [59] 37B 8.8 38.9 18.2 0.0 0.0
VILA1.0-13b [65] 13B 15.7 41.9 33.2 8.6 0.1
llava-v1.6-13b [23] 13B 16.9 43.7 34.9 13.6 0.0
llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
InternVL1.5-RAG [11] 45B 46.1 59.5 50.1 44.9 39.3
@@
Table 8: Performance comparison with other DocVLMs on MM-NIAH. The evaluation approach aligns with the benchmark.
@@ 5.4 Ablation Study
To comprehensively evaluate the effectiveness of our contributions, we conduct ablation studies focusing on the impact of the sparse sampler, the dataset, the document length, and sampling strategies. These experiments are based on English PaperPDF for training and evaluation to avoid interference from other factors.
@@
Table 9: Ablation study on the impact of sparse sampler
@@
We retrain PDF-WuKong on various subsets of our English PaperPDF, and the results are shown in Tab. 10. Increasing the amount of training data leads to consistent improvements in the model’s accuracy, proving that our dataset follows scaling laws. In addition, we verify the effectiveness of our two data construction methods. Under the same data scale, multi-evidence data can enhance the model’s complex reasoning ability.
@@
To understand the impact of document length on model performance and efficiency, we divide the test set into subsets based on the number of pages per document. Results in Fig. 4 demonstrate that our model’s performance and token counts remain relatively stable across documents of varying lengths. This stability indicates that the sparse sampler effectively reduces the input size to a reasonable level, regardless of the original document length. In contrast, the baseline MLLM without the sparse sampler is unable to handle long documents effectively. Its performance deteriorates significantly as the document length increases. The number of tokens also increased dramatically, resulting in huge resource consumption. These findings highlight the robustness of our model in processing long documents without sacrificing accuracy or incurring extra computational costs.
@@
We explore the impact of different numbers of text blocks or diagrams selected by the sparse sampler. As shown in Tab. 11, setting a small top $k$ can lead to missing crucial information needed for accurately answering queries, thus reducing performance. Conversely, a larger top $k$ introduces redundant information and increases computational costs without significantly enhancing accuracy. To strike a balance between performance and resource efficiency, we use the top 5 as the default setting.
Sampling chunks ANLS F1 ROUGE Tokens
Top 1 39.20 38.28 35.26 1186
Top 3 42.09 43.06 39.69 1452
Top 5 42.59 43.63 40.19 1789
Top 10 43.01 44.22 40.67 2386
Top 15 43.19 44.57 42.08 2704
Top 20 43.42 45.02 42.30 3364
Table 11: Ablation study of different sampling strategy
@@ 6 Conclusion
Appendix A Algorithm
Algorithm 1 shows the detailed inference process of PDF-WuKong. The training pipeline is shown in Algorithm 2. Our PDF-WuKong can achieve efficient and accurate understanding of long PDFs with end-to-end sparse sampling.
Algorithm 1 Inference pipeline for PDF-WuKong
1:Input: PDF document $D$, user query $q$
2:Output: Generated answer $a$
3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
4:Stage 1: Document Parsing
5:Parse the input document $D$ into text blocks and images:
$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
6:Stage 2: Sparse Sampling
$E_{T}=\{e_{T_{1}},e_{T_{2}},\dots,e_{T_{n}}\}\leftarrow\text{En\_T}(\{T_{1},T_{2},\dots,T_{n}\}),$
$E_{I}=\{e_{I_{1}},e_{I_{2}},\dots,e_{I_{m}}\}\leftarrow\text{En\_I}(\{I_{1},I_{2},\dots,I_{m}\})$
8:Encode the user query $q$:
$e_{q}\leftarrow\text{En\_T}(q)$
9:Calculate the similarity between query embedding $e_{q}$ and cached text/image embeddings $\{E_{T},E_{I}\}$:
$S_{T}=\{\text{Sim}(e_{q},e_{T_{i}})\mid i=1,2,\dots,n\},$
$S_{I}=\{\text{Sim}(e_{q},e_{I_{j}})\mid j=1,2,\dots,m\}$
10:Select the top-$k$ relevant text blocks and images:
$(T,I)_{top}\leftarrow\text{TopK}(S_{T},S_{I},k)$
11:Stage 3: Answer Generation
12:Input the query $q$ and the selected tokens into the LLM:
$a\leftarrow\text{LLM}(q,(T,E_{I})_{top})$
13:Return the generated answer $a$.
Appendix B Visualization Results
B.1 Qualitative Comparison with Other Products
The qualitative comparison between PDF-WuKong and other proprietary products on the long PDF understanding task is presented in Fig. 5. PDF-WuKong achieves significant performance advantages over others such as Qwen [66], ChatGLM [67], Kimi [68], and Gemini-pro [69]. The success stems from its ability to effectively sample key evidence relevant to the question and its strong multimodal understanding capabilities. This observation further validates the effectiveness of our method in enhancing the long PDF understanding capabilities of MLLMs.
Algorithm 2 Training pipeline for PDF-WuKong
1:Input: PDF document $D$, user query $q$, ground truth answer $gt$
2:Output: Final loss function $\mathcal{L}_{\text{total}}$
3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
4:Stage 1: Data Preparing
5:Text blocks and images:
$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
6:Stage 2: Multimodal encoding
7:Encode the user query, positive and negative samples:
$e_{q}\leftarrow\text{En\_T}(q)$
$E_{T}=\{e_{T_{P}},e_{T_{N}}\}\leftarrow\text{En\_T}(\{T_{P},T_{N}\}),$
$E_{I}=\{e_{I_{P}},e_{I_{N}}\}\leftarrow\text{En\_I}(\{I_{P},I_{N}\}),$
8:Calculate the contrastive learning loss:
$\mathcal{L}_{\text{rep}}(e_{q},\{e_{T_{P}},e_{I_{P}}\},\{e_{T_{N}},e_{I_{N}}\})$
9:Stage 3: Output prediction of MLLM
10:Input the query, the positive text, and the positive image tokens from the shared image encoder En_I:
$a\leftarrow\text{LLM}(q,T_{P},e_{I_{P}})$
11:Calculate the cross-entropy loss:
$\mathcal{L}_{\text{QA}}(a,gt)$
12:Stage 4: Optimize model in an end-to-end manner
13:Update model parameters according to the joint loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}$
14:Return the final loss $\mathcal{L}_{\text{total}}$.
Figure 5: Qualitative comparison between PDF-WuKong with other proprietary products. The red box indicates the evidence that the correct answer depends on.
B.2 Visualization Results on Chinese Documents
In addition to the outstanding performance on English documents, PDF-WuKong also demonstrates remarkable capabilities on Chinese documents. Fig. 6 shows its qualitative performance on the Chinese long PDF document understanding task. It effectively samples relevant evidence within the PDF and demonstrates the ability to derive accurate answers based on the sampled evidence.
Appendix C More Dataset Details
The Q-E-A triplets in the PaperPDF dataset are derived from approximately 60k English and 10k Chinese documents, encompassing nearly 70 disciplines such as computer science, engineering, and materials science.
To comprehensively illustrate the characteristics of the PaperPDF dataset, we conduct a detailed statistical analysis. The key statistics of the source documents in PaperPDF are presented in Tab. 12. Compared to previous multi-page document datasets, PaperPDF not only includes a significantly larger number of documents but also features more pages and OCR tokens. More importantly, PaperPDF encompasses both Chinese and English documents with interleaved text and images. Our dataset provides an important innovation driver and a comprehensive benchmark for the development of the multimodal long PDF document understanding task.
@@ Appendix C More Dataset Details
For the five types of data (Text-only, Image-only, Image-text, Section, Cross-paragraph) in single evidence and multiple evidence, Fig. 7 - Fig. 11 show their prompt engineering and a corresponding example, respectively.
MPdataset Document Language Page Image Token
DUDE [17] 5k En 6 - 1831
MP-Docvqa [46] 6k En 8 - 2026
PaperPDF (En) 60k En 25 12 11371
PaperPDF (Zh) 10k Zh 11 7 3413
PaperPDF 70k En+Zh 23 11 10234
Table 12: Statistics of the source documents, where Page, Image, and Token represent the average number of pages per document, the average number of images in the document, and the average number of OCR tokens per document, respectively.
@@
Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure? Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1). Figure: Table 1[Uncaptioned image] Caption: Experimental results of ablation analysis, different training strategies of our method and comparison with other approaches. Note: the -means the results were not reported by that method. Answer 1: Ours. Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks. This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
Appendix D Limitation

### v3

@@ 5.3 Document-oriented VQA
To validate the strong capability of our model in other document understanding scenarios, we conduct experiments on several public benchmarks. It is worth noting that while our primary pipeline utilizes parsed PDF content, PDF-WuKong is natively compatible with non-PDF inputs. In the following experiments, we feed the model with page images rather than parsed PDF blocks. PDF-WuKong treats each page as a visual block and can directly retrieve the pages containing the evidence.
First, we evaluate the performance of PDF-WuKong on single-page document datasets Mathew_2021_WACV ; masry2022chartqa ; Mathew_2022_WACV . As shown in Tab. 8, our model achieves comparable performance on single-page document understanding. This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question-answering tasks.
@@
Qwen-VL Qwen-VL 65.1 65.7 35.4
Monkey li2023monkey 66.5 65.1 36.1
Text-Monkey liu2024textmonkey 73.0 66.9 28.6
MiniCPM-V-2.5 hu2024minicpm 84.8 - -
Vary-base wei2025vary 76.3 66.1 -
TextHawk yu2024texthawk 76.4 66.6 50.6
IXC2-4KHD-16 internlmxcomposer2_4khd 84.9 80.1 60.8
DocOwl 2 hu2024mplugdocowl2 80.7 70.0 46.4
CREAM zhang2024cream 79.4 - 53.6
@@
LayoutLMv3 huang2022layoutlmv3 55.1 20.3
Longformer beltagy2020longformer 55.1 27.1
BigBird zaheer2020big 58.5 26.3
Hi-VT5 TITO2023109834 61.8 35.7
DocFormerv2 appalaraju2024docformerv2 76.4 48.4
GRAM blau2024gram 83.0 53.4
GPT-4V (2024-06) 2023gpt4V - 53.9
Idefics3-8B laurenccon2024obelics 67.2 38.7
DocOwl2 hu2024mplugdocowl2 69.4 46.7
CREAM zhang2024cream 65.3 52.5
SV-RAG chen2025svrag 71.0 45.0
@@
In addition, we assess the performance of specialized models and MLLMs on two existing multi-page document QA datasets. The results shown in Tab. 9 indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document MLLM DocOwl2 hu2024mplugdocowl2 . Notably, on complex multi-page document datasets like DUDE van2023document , PDF-WuKong outperforms GPT-4V 2023gpt4V . This improvement is attributed to the sparse sampler, which extracts useful information from multi-page documents, enabling the model to focus on relevant content.
Furthermore, we conduct zero-shot evaluations on a long multimodal document understanding benchmark MM-NIAH wang2024needle . As shown in Tab. 10, our model uses the fewest parameters yet achieves the second-best performance. Although InternVL-V1.5-RAG surpasses PDF-WuKong by 2.8%, it utilizes 36.5B more parameters than our model. Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, while that of other models significantly decreases. At a context length of 64K, PDF-WuKong even achieves the best performance, demonstrating its robustness in handling long-context multimodal inputs.
Table 10: Performance comparison with other DocVLMs on $M$-NIAH. The evaluation approach aligns with the benchmark.
@@
Emu2-Chat sun2024generative 37B 8.8 38.9 18.2 0.0 0.0
VILA1.0-13b lin2024vila 13B 15.7 41.9 33.2 8.6 0.1
llava-v1.6-13b liu2024llavanext 13B 16.9 43.7 34.9 13.6 0.0
llava-v1.6-34b liu2024llavanext 34B 20.6 57.4 45.1 8.2 0.0
InternVL1.5 chen2024far 26B 41.1 59.5 50.1 41.9 16.6
InternVL1.5-RAG chen2024far 45B 46.1 59.5 50.1 44.9 39.3
@@
5.4 Fine-grained Evaluation on Long PDF Understanding
To comprehensively assess the capabilities of PDF-WuKong in long-document processing, we conducted a fine-grained evaluation on the PaperPDF benchmark, as presented in Table 11. We can observe that the model exhibits superior performance in figure and table understanding tasks compared to other categories questions, validating that the sparse sampling mechanism effectively captures critical visual evidence within long documents. However, formula understanding remains weaker, which may be attributed to the loss of 2D structural semantics during text linearization and the scarcity of relevant training data. Additionally, the model performs slightly better on Multi-hop tasks, a trend suggesting that the model potentially benefits from utilizing global redundant context, unlike single-hop tasks that focus on locating isolated details.
Table 11: Performance of PDF-WuKong across different question types. TU, FU, FoU, MuU, MuL, SiU and SiL represent Table Understanding, Figure Understanding, Formula Understanding, Multi-hop Understanding, Multi-hop Logical Reasoning, Single-hop Understanding and Single-hop Logical Reasoning, respectively.
Question PDF-WuKong InternVL2
Type ANLS F1 ROUGE ANLS F1 ROUGE
TU 44.15 46.32 45.09 40.65 45.44 45.71
FU 46.05 48.38 46.46 40.82 44.68 45.03
FoU 39.67 39.81 38.08 26.67 28.59 30.72
MuU 44.07 45.30 45.01 35.83 37.07 39.13
MuL 41.13 42.40 38.92 32.13 33.37 34.26
SiU 42.32 43.33 40.30 28.69 31.11 30.90
SiL 39.05 38.41 35.55 27.04 28.16 28.46
@@ 5.5 Ablation Study
To comprehensively evaluate the effectiveness of our contributions, we conduct ablation studies focusing on the impact of the sparse sampler, the dataset, the document length, and sampling strategies. These experiments are based on the English subset of our PaperPDF for training and evaluation to avoid interference from other factors.
@@
Table 12: Ablation study on the impact of sparse sampler
@@
We retrain PDF-WuKong on various dataset settings of the English subset of our PaperPDF, and the results are shown in Tab. 13. Increasing the amount of training data leads to consistent improvements, proving that our dataset is of high quality and follows scaling laws. Besides, we verify the effectiveness of our two data construction methods. Under the same data scale, multi-evidence data can enhance the model’s complex reasoning ability.
Table 13: Ablation study on dataset setting
@@
To understand the impact of document length on model performance and efficiency, we divide the test set into subsets based on the number of pages per document. Results in Fig. 7 show that our model’s performance and token counts remain stable across documents of varying lengths. This stability indicates that the sparse sampler effectively reduces the input size, regardless of the original document length. In contrast, the baseline without the sparse sampler is unable to handle long documents effectively. Its performance deteriorates significantly as the document length increases. The number of tokens also increased dramatically, resulting in huge resource consumption. These findings highlight the robustness of our model in processing long documents without sacrificing accuracy or incurring extra computational costs.
@@
Table 14: Ablation study of different sampling strategy. P and R, represent the Precision and Recall of the evidence sampling module.
Sampling Retrieval QA Token
chunks P R ANLS F1 ROUGE
Top 1 46.00 45.60 39.20 38.28 35.26 1186
Top 3 22.53 66.72 42.09 43.06 39.69 1452
Top 5 14.96 73.46 42.59 43.63 40.19 1789
Top 10 8.40 81.79 43.01 44.22 40.67 2386
Top 15 5.94 86.61 43.19 44.57 42.08 2704
Top 20 4.63 89.64 43.42 45.02 42.30 3364
@@
To investigate the effectiveness of the sparse sampler in evidence location, we conduct an ablation study on the number of selected blocks. Beyond standard QA performance evaluation, we further explore the precision and recall of the sampled evidence , following the previous studies souibgui2025docvxqa ; chen2025towards . As shown in Tab. 14, setting a small top $k$ results in relatively low recall, which implies that the crucial evidence required for accurately answering queries is often omitted, thereby constraining the model’s performance. Conversely, although a larger top-$k$ indeed improves evidence recall, it comes at the cost of a sharp decline in precision, introducing substantial non-evidential redundancy that interferes with the model’s reasoning. As a result, when top-$k$ exceeds 5, the model exhibits diminishing returns, where a substantial increase in token consumption yields only marginal gains in QA accuracy. To strike an optimal balance between performance and resource efficiency, we adopt Top-5 as the default setting.
@@ 6 Conclusion
7 Limitation
It is important to note that unlike the manually verified test set, the training set is constructed via an automated pipeline. Consequently, it may contain a certain degree of noise stemming from PDF parsing errors (e.g., imperfect layout analysis in complex documents) or alignment inconsistencies in multi-evidence extraction. While this noise is typical in large-scale pre-training datasets, users should be aware of these potential imperfections.
Data Availability Statement
The datasets used during and/or analysed in the current study are available in our PaperPDF repository [Link], the DocVQA repository [Link], the ChartQA repository [Link], the InfoVQA repository [Link], the MP-DocVQA repository [Link], the DUDE repository [Link], and the $M$-NIAH repository [Link].
Declaration
The authors have no relevant financial or non-financial interests to disclose.
Acknowledgement
This research is supported by the National Science Fund for Distinguished Young Scholars of China (Grant No.62225603) and the National Natural Science Foundation of China (Grant No. 62576147).
Appendix A Prompts and Examples for PaperPDF Dataset
@@
Fig. 8 - Fig. 12 show the prompt engineering for constructing the five types of data (Text-only, Image-only, Image-text, Section, Cross-paragraph) in single evidence and multiple evidence. There are also corresponding examples for these five types.